OpenStudy (nuccioreggie):

How many solutions does the equation 4s + 8 = 3 + 5 + 4s have? Two None One Infinitely many

OpenStudy (whovianchick):

4s+8=3+5+4s 4s+8= 8+4s 0=0 No solutions (Answer B)

OpenStudy (whovianchick):

Since you can't solve for s, it just cancels out and there are no solutions. Get it?

OpenStudy (nuccioreggie):

ok your making alot of since and im understanding better

OpenStudy (mathmale):

This is a linear equation, so the most solution(s) it could have is ONE. But it's possible that there's no solution. Please do the actual work of solving for s, as "whovianchick" has done.

OpenStudy (whovianchick):

That's good :)

OpenStudy (landon34):

you can try out the equation solver. URL: http://www.geteasysolution.com/

OpenStudy (whovianchick):

Oh yeah @mathmale made a good point if nothing is squared, then there won't be 2 solutions.